formulae for motion of bodies in motion

Grade:11

2 Answers

Gayatri Jayesh Bondriya
8 years ago
First law: When viewed in an internal refrance  frame , an object either remains at rest or continues to move at a constant velocity  unless acted upon by an external force
Second law: The vector  sum of the external forces F on an object is equal to the mass m of that object multiplied by the accleration  vector a of the object: F = ma.
Third law: When one body exerts a force on a second body, the second body simultaneously exerts a force equal in magnitude and opposite in direction on the first body.
Gayatri Jayesh Bondriya
8 years ago
The first law states that if the net force  (the vector  sum  of all forces acting on an object) is zero, then the velocity  of the object is constant.
 

The second law states that the net force  on an object is equal to the rate of change of its linear   momentomm  p in an internal refrance  frame

\mathbf{F} = \frac{\mathrm{d}\mathbf{p}}{\mathrm{d}t} = \frac{\mathrm{d}(m\mathbf v)}{\mathrm{d}t}.

The second law can also be stated in terms of an object's acceleration. Since Newton's second law is only valid for constant-mass systems mass can be taken outside the diffrenciation operator . Thus,

BF = −AF, and the two forces are equal and opposite: A on BF simultaneously exerts a force B, then B on a second object AF exerts a force AThe third law states that all forces between two objects exist in equal magnitude and opposite direction: if one object   \mathbf{F} = m\,\frac{\mathrm{d}\mathbf{v}}{\mathrm{d}t} = m\mathbf{a},